OpenStudy (anonymous):

Which set of points are on the line y = 8x - 6?

OpenStudy (anonymous):

you cant substitude 0 for x and y respectively

OpenStudy (anonymous):

(0)=8x-6 6=8x x=6/8 x=3/4 (3/4,0)

OpenStudy (anonymous):

do the same for other variable